OpenStudy (anonymous):

Find the value of y when x equals 1 in this equation. 2x + (-7y) = -54 A. 8 B. 6 C. -8 D. -6

OpenStudy (anonymous):

2 + (-7y)= -54 -7y = -56 y = 8

OpenStudy (umarsback):